# A Survey on Deep Learning based Time Series Analysis with Frequency Transformation

### v5

@@ 7.1.2. Integrating Frequency Transformation with Deep Learning
The basis functions used in frequency transformation, such as sine, cosine, and wavelet functions, are fixed across different domains. As a result, the frequency features extracted through these basis functions are domain-invariant. In other words, the features are insensitive to unexpected noise or to changing conditions. Few previous works combine frequency transformation with the learning ability of neural networks. (Khan_nips_2018, ) proposes a method to efficiently optimize the parameters of the spectral decomposition based on the wavelet transform in a neural network framework. (pnas_2022, ) mimics the fast DWT cascade architecture utilizing the deep learning framework. These methods have shown promising performances, and in the future, the combination of frequency transformation with deep learning deserves further investigation.
@@ 7.1.3. Jointly Learning in the Time and Frequency Domain
The frequency domain only uses periodic components, and thus cannot accurately model the non-periodic aspects of a signal, such as a linear trend (GodfreyG_tnnls_2018, ). Moreover, according to the uncertainty principle (TFAD_2022, ), designing a model with a single structure that can capture the time and frequency patterns simultaneously is difficult. As a result, in the future, an interesting direction is to take advantage of corresponding characteristics of learning in the time and frequency domain to improve the accuracy and efficiency of time series analysis. Few works have tried to learn representations in the time and frequency domain, respectively. More time-frequency representation learning methods are required in the future.
@@ 7.2.2. Scalability
Scalability (KeoghK03, ) is a key consideration in time series analysis. When coupled with frequency transformation techniques, it offers the potential for efficient and scalable analysis of large-scale time series data. Frequency transformation allows for the extraction of frequency components, reducing the dimensionality of the data and enabling more efficient processing. This dimensionality reduction can significantly improve the scalability of time series analysis algorithms, as it reduces computational complexity and memory requirements. Scalable time series analysis with frequency transformation can pave the way for analyzing and extracting insights from big data time series applications in domains such as the Internet of Things (IoT), financial markets, or sensor networks.
@@ 7.2.3. Privacy-preserving
Leveraging frequency transformation offers a powerful approach to data privacy-preserving (DworkMNS16, ) in time series analysis. By applying frequency transformation, time series data can be transformed into frequency domain representations without revealing the underlying raw data. This transformation allows for the extraction of frequency components and patterns while maintaining the confidentiality of the original information. Privacy-preserving with frequency transformation techniques can ensure individual privacy and data confidentiality, and enable collaborative analysis, data sharing, and research collaborations while mitigating privacy risks. This approach is particularly valuable in domains where data sensitivity is critical, such as healthcare, finance, or personal monitoring, allowing for the utilization of frequency analysis while protecting the privacy of individuals or organizations.
@@ 8. Conclusion
In this paper, we provide a comprehensive survey on deep learning based time series analysis with frequency transformation. We organize the reviewed methods from the perspectives of incorporation approaches, neural network design, and application domains, and we summarize the advantages and limitations of frequency transformation for time series analysis. To the best of our knowledge, this paper is the first work to comprehensively and systematically review neural time series analysis with frequency transformation, which would greatly benefit the time series community. Additionally, we offer a curated collection of sources, accessible at https://github.com/aikunyi/time_series_frequency, to further assist the research community.

### v6

@@ 7.1.2. Integrating Frequency Transformation with Deep Learning
The basis functions used in frequency transformation, such as sine, cosine, and wavelet functions, are fixed across different domains. As a result, the frequency features extracted through these basis functions are domain-invariant. In other words, the features are insensitive to unexpected noise or to changing conditions. Few previous works combine frequency transformation with the learning ability of neural networks. (Khan and Yener, 2018) proposes a method to efficiently optimize the parameters of the spectral decomposition based on the wavelet transform in a neural network framework. (Michau et al., 2022) mimics the fast DWT cascade architecture utilizing the deep learning framework. These methods have shown promising performances, and in the future, the combination of frequency transformation with deep learning deserves further investigation.
@@ 7.1.3. Jointly Learning in the Time and Frequency Domain
The frequency domain only uses periodic components, and thus cannot accurately model the non-periodic aspects of a signal, such as a linear trend (Godfrey and Gashler, 2018). Moreover, according to the uncertainty principle (Zhang et al., 2022b), designing a model with a single structure that can capture the time and frequency patterns simultaneously is difficult. As a result, in the future, an interesting direction is to take advantage of corresponding characteristics of learning in the time and frequency domain to improve the accuracy and efficiency of time series analysis. Few works have tried to learn representations in the time and frequency domain, respectively. More time-frequency representation learning methods are required in the future.
@@ 7.2.2. Scalability
Scalability (Keogh and Kasetty, 2003) is a key consideration in time series analysis. When coupled with frequency transformation techniques, it offers the potential for efficient and scalable analysis of large-scale time series data. Frequency transformation allows for the extraction of frequency components, reducing the dimensionality of the data and enabling more efficient processing. This dimensionality reduction can significantly improve the scalability of time series analysis algorithms, as it reduces computational complexity and memory requirements. Scalable time series analysis with frequency transformation can pave the way for analyzing and extracting insights from big data time series applications in domains such as the Internet of Things (IoT), financial markets, or sensor networks.
@@ 7.2.3. Privacy-preserving
Leveraging frequency transformation offers a powerful approach to data privacy-preserving (Dwork et al., 2016) in time series analysis. By applying frequency transformation, time series data can be transformed into frequency domain representations without revealing the underlying raw data. This transformation allows for the extraction of frequency components and patterns while maintaining the confidentiality of the original information. Privacy-preserving with frequency transformation techniques can ensure individual privacy and data confidentiality, and enable collaborative analysis, data sharing, and research collaborations while mitigating privacy risks. This approach is particularly valuable in domains where data sensitivity is critical, such as healthcare, finance, or personal monitoring, allowing for the utilization of frequency analysis while protecting the privacy of individuals or organizations.
@@ 8. Conclusion
In this paper, we present a comprehensive survey on deep learning based time series analysis methods that leverage frequency transformation. We organize the reviewed methods from the perspectives of incorporation approaches, neural network design, and application domains, and we summarize the advantages and limitations of frequency transformation for time series analysis. To the best of our knowledge, this is the first systematic and in-depth review focused specifically on neural time series analysis with frequency transformation, aiming to provide valuable insights for the research community. To support further exploration, we also provide a curated repository of related resources, available at https://github.com/aikunyi/time_series_frequency.
Acknowledgements.
This project is funded by the State Grid Corporation of China headquarters science and technology project ”Research on the optimization technology and platform of the company’s natural monopoly business and competitive business operation based on social information” (Project No. 52180025000C-194-ZN).